OpenStudy (anonymous):

Sarosh receives 8% commission on the final price of the clothing he sells. A customer bought a $45 coat that was marked 20% off and a $95 suit that was marked 30% off. What would Sarosh's commission be on that customer's purchase?

OpenStudy (anonymous):

First you want to find the total price of the coat and suit (after they were marked down) The coat was marked 20% off of 45, so we do 45*0.20 = 9 So the coat was marked down $9, we subtract this from 45 to see what the coat actually sold for. 45-9 = 36 Now we do the same thing for the suit. It was 30% off of 95. 95*0.30 = 28.5 95-28.5 =66.50 The final sales price is $66.50. See if you can find 8% of 66.5 to get your final answer!

OpenStudy (anonymous):

Thank you so much !